Moonshine has 50 liters of 70% alcoholic solution. How many liters of pure alcohol would you need to add to make an 80% solution?
70% solution 50liters .70(50 + p)
Alcohol p 1.00p
80% solution 50 + p .80(50 + p)

?

Answer by Paul(988)

0.8(50-t)=0.7(50)
40-0.8t=35
-0.8t=-5
t=6.25
Therefore about 6.25 liters of alcohol would be needed to make 80% of the soulution